11. Why is it a good idea to buy two lottery tickets as opposed to one lottery ticket if you want to hit the jackpot? Explain the rationale using the logic of probability, in particular using the language of AND/OR. Which of these (AND/OR) is appropriate in explaining the increased odds.
Answer #1

it is a good idea to buy two lottery tickets as opposed to one lottery ticket if you want to hit the jackpot.

because probability of winning becomes twice.

let P(win with one ticket) = p

then , P(winning with two tickets) = P(win one ticket OR win second ticket)

= P(winning first  ticket) + P(winning second tickets) = p + p = 2p

hence, probability of winning gets double

here, OR will be appropriate to use